stylequeen:

Choose the equation below that represents the line that passes through the point (7, −2) and has a slope of −3. (5 points) y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2)

supie:

Is this a test question?

Timmyspu:

ok so what do you know is not going to be the correct answer

stylequeen:

3 and 4

darkknight:

To write this in point-slope form, we need to know the slope and 1 point, which we have \[y-y1=m(x-x1)\] \[y+2=-3(x-7)\] Because we plug in -2 for y1 7 for x1 and -3 for m

stylequeen:

ok thank u

darkknight:

np